# Shuffle Products, Degenerate Affine Hecke Algebras, and Quantum Toda Lattice

### v2

@@ 7.2. Monopole operators
$S^{k}\mathbf{C}^{N}\ni v_{1}^{\otimes k}\mapsto\left(\sum_{i=1}^{N}\prod_{j\neq i}\frac{1}{w_{i}-w_{j}}\mathbf{u}_{i}^{-1}\right)^{k}=:\sum_{J}c_{J}(w)\mathbf{u}_{j_{1}}^{-1}\ldots\mathbf{u}_{j_{k}}^{-1},$
where $J$ runs over all multisubsets of $\{1,\ldots,N\}$ and $c_{J}(w)\in\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}$ are some functions. In what follows, we will use Proposition˜4.12 and notations thereof. By Proposition˜7.16, we need to compute
$\sum_{J}c_{J}(w)f(\vec{Y})\mathbf{u}^{-1}_{j_{1}}\ldots\mathbf{u}^{-1}_{j_{k}},\qquad f(\vec{Y})=\frac{1}{k!}\prod_{i\neq j}(Y_{i}-Y_{j}-\hbar).$
@@
Therefore, the only non-zero indices in the sum are permutations of $\{1,\ldots,k\}$. Moreover, by $W$-invariance of Remark˜7.10, it is enough to compute the coefficient $c_{\{1,\ldots,k\}}(w)$. By an easy induction on $k$, we have
$c_{\{1,\ldots,k\}}(w)=k!\prod_{1\leq i\neq j\leq k}\frac{1}{w_{i}-w_{j}-\hbar}\cdot\prod_{\begin{subarray}{c}i=1\ldots k\\ j=k+1\ldots N\end{subarray}}\frac{1}{w_{i}-w_{j}}\mathbf{u}^{-1}_{1}\ldots\mathbf{u}^{-1}_{k}.$
By Proposition˜4.12,
$c_{\{1,\ldots,k\}}(w)f(\vec{Y})\mathbf{u}_{1}^{-1}\ldots\mathbf{u}_{k}^{-1}=\prod_{\begin{subarray}{c}i=1\ldots k\\ j=k+1\ldots N\end{subarray}}\frac{1}{w_{i}-w_{j}}\mathbf{u}_{1}^{-1}\ldots\mathbf{u}_{k}^{-1}.$
@@
$\langle g\cdot v_{\omega_{k}},v^{*}_{\omega_{k}}\rangle\mapsto\sum_{\begin{subarray}{c}I\subset\{1,\ldots,N\}\\ |I|=k\}\end{subarray}}\prod_{\begin{subarray}{c}i\in I\\ j\in\bar{I}\end{subarray}}\frac{1}{w_{i}-w_{j}}\prod_{i\in I}\mathbf{u}_{i}^{-1},$
@@
This is a rational analog of the simplest operator in the M-system in its shuffle realization, see [Tsy23, (104)] and [Tsy23, Proposition 6.13]. By [FT19, Remark 9.19], its operators correspond to the $K$-theory classes of line bundles $\mathcal{O}(-k)$ restricted to the orbits $\mathrm{Gr}^{\omega_{i}}$ as in Section˜7.1, where $\mathcal{O}(1)$ is the theta bundle on the affine Grassmannian. Therefore, Proposition˜7.16 and Theorem˜7.17 are in accordance with the derived Satake isomorphism of Theorem˜7.14, since by [Fin+18, Lemma 2.6], we have $\langle g\cdot v_{\omega_{i}},v_{\omega_{i}}^{*}\rangle\mapsto[\mathrm{Gr}^{\omega_{i}}]$, up to a scalar factor. It would be interesting to find a rational analog of the full $M$-system.
@@ Appendix A Technical lemmas
The following result is used in the proof of Proposition˜2.3. Recall that
@@
$\prod_{i=1}^{l}G^{\theta}(u-z_{i},\kappa)=\sum_{i=1}^{l}G^{\theta}(u-z_{i},l\kappa)\cdot\prod_{\begin{subarray}{c}j=1\ldots l\\ j\neq i\end{subarray}}G^{\theta}(z_{i}-z_{j},\kappa).$
@@
$\frac{\theta(u-z_{i}+l\kappa)\theta(u-z_{l+1}+\hbar)}{\theta(u-z_{i})\theta(l\kappa)\theta(u-z_{l+1})\theta(\kappa)}=$
$=$ $\frac{\theta(u-z_{l+1}+l\kappa+\kappa)\theta(z_{l+1}-z_{i}+l\kappa)}{\theta(u-z_{l+1})\theta(l\kappa+\kappa)\theta(z_{l+1}-z_{i})\theta(l\kappa)}+\frac{\theta(z_{i}-z_{l+1}+\kappa)\theta(u-z_{i}+l\kappa+\kappa)}{\theta(z_{i}-z_{l+1})\theta(\kappa)\theta(u-z_{i})\theta(l\kappa+\kappa)}$
by mutiplying by common denominator and taking
@@
The following result about symmetric polynomials is used in Proposition˜5.7. Define $p^{\hbar}_{a}(\vec{\Omega})\in\mathbf{C}[\hbar][\Omega_{1},\ldots,\Omega_{k}]^{S_{k}}$ by
$\prod_{i=1}^{k}\frac{u-\Omega_{i}+\hbar}{u-\Omega_{i}}=1+\hbar\sum_{a=0}^{\infty}p_{a}^{\hbar}(\vec{\Omega})u^{-a-1}\in\mathbf{C}[\hbar][\Omega_{1},\ldots,\Omega_{k}]^{S_{k}}[[u^{-1}]],$
where we understand each term in the product as in Remark˜4.8.
@@
    (A.1) $p_{a}^{\hbar}(\vec{\Omega})=\sum_{i=1}^{k}\Omega_{i}^{a}\prod_{j\neq i}\frac{\Omega_{i}-\Omega_{j}+\hbar}{\Omega_{i}-\Omega_{j}}.$
    In particular, $p^{0}_{a}(\vec{\Omega})=\sum_{i=1}^{k}\Omega_{i}^{a}$ is the standard power sum symmetric function.
  2. (2)
    The functions $p^{\hbar}_{a}(\vec{\Omega})$ for $1\leq a\leq k$ freely generate the algebra of symmetric functions, in other words, there is an isomorphism
    $\mathbf{C}[\hbar][\Omega_{1},\ldots,\Omega_{k}]^{S_{k}}\cong\mathbf{C}[\hbar][p^{\hbar}_{1}(\vec{\Omega}),\ldots,p^{\hbar}_{k}(\vec{\Omega})].$
@@
$\mathbf{C}[\hbar][\Omega_{1},\ldots,\Omega_{k}]^{S_{k}}\cong\mathbf{C}[\hbar][p^{0}_{1}(\vec{\Omega}),\ldots,p^{0}_{k}(\vec{\Omega})],$
since $p^{0}_{a}(\vec{\Omega})=\sum_{i=1}^{k}\Omega_{i}^{a}$. Let us consider the right-hand side as the coordinate ring of the affine space $\mathbb{A}^{k}\times\mathbb{A}^{1}$, where $\mathbb{A}^{1}$ corresponds to $\hbar$. There is a $\mathbf{C}^{*}$-action on $\mathbf{C}[\hbar][\Omega_{1},\ldots,\Omega_{k}]$ such that all the variables $\Omega_{i}$ and $\hbar$ have weight 1. It induces a $\mathbf{C}^{*}$-action on $\mathbf{C}[\hbar][\Omega_{1},\ldots,\Omega_{k}]^{S_{k}}$ with positive weights; in particular, it contracts $\mathbb{A}^{k}\times\mathbb{A}^{1}$ to the origin. In fact, under the isomorphism above, this action is linear and diagonal with weights $(1,\ldots,k,1)$.
Consider the map $\varphi_{\hbar}\colon\mathbb{A}^{k}\times\mathbb{A}^{1}\rightarrow\mathbb{A}^{k}\times\mathbb{A}^{1}$ defined on the ring of functions by $p^{0}_{a}(\vec{\Omega})\mapsto p^{\hbar}_{a}(\vec{\Omega})$ for every $a$ and $\hbar\mapsto\hbar$. It follows from (A.1) that this map is $\mathbf{C}^{*}$-equivariant. Moreover, it is clear from weight considerations and explicit calculations for $a=1$ that the linear part of $\varphi_{\hbar}$ is
$\mathrm{d}\varphi_{\hbar}(p^{0}_{a}(\vec{\Omega}))=p_{a}^{0}(\vec{\Omega})+\delta_{a1}\frac{k(k-1)}{2}\hbar,\qquad\mathrm{d}\varphi_{\hbar}(\hbar)=\hbar.$
Therefore, the differential of $\varphi_{\hbar}$ at zero is invertible. By the proof of [GG02, Lemma 2.1], it is an isomorphism. In particular, the polynomials $\{p^{\hbar}_{a}(\vec{\Omega})\}$ generate the ring of symmetric functions. ∎
In what follows, we use the notations of Section˜4 for $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes(S^{\bullet}\mathbf{C}^{N}\otimes S^{\bullet}(\mathbf{C}^{N})^{*})$ (more precisely, for its Mickelsson algebra).
@@
$P\sum_{1\leq i_{1}<\ldots<i_{s}<i}(w_{i}-w_{i_{1}})^{-1}\ldots(w_{i}-w_{i_{s}})^{-1}\phi_{i_{1}}E_{i_{1}i_{2}}\ldots E_{i_{s}i}\cdot v_{j}P=$
$=$ $Pv_{j}\sum_{1\leq i_{1}<\ldots<i_{s}<i}(w_{i}-w_{i_{1}})^{-1}\ldots(w_{i}-w_{i_{s}})^{-1}\phi_{i_{1}}E_{i_{1}i_{2}}\ldots E_{i_{s}i}P=Pv_{j}\phi_{i}P$
@@
$P\sum_{N\geq j_{1}>\ldots>j_{s}>j}(w_{j}-w_{j_{1}})^{-1}\ldots(w_{i}-w_{j_{s}})^{-1}v_{j_{1}}E_{j_{2}j_{1}}\ldots E_{jj_{s}}E_{i_{1}i_{2}}\cdot\phi_{i}P=P\phi_{i}v_{j}P.$
@@
$\varpi_{N}\colon\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes(S^{\bullet}\mathbf{C}^{N}\otimes S^{\bullet}(\mathbf{C}^{N})^{*})\rightarrow[\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes(S^{\bullet}\mathbf{C}^{N}\otimes S^{\bullet}(\mathbf{C}^{N})^{*})]^{\mathrm{opp}},\qquad\varpi_{N}(v_{i})=\phi_{i},\qquad\varpi_{N}(\phi_{i})=v_{i},$
extending the one (1.6) of $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})$. By Remark˜4.3 we have $\varpi_{N}(P)=P$, therefore,
@@
The proof of the following proposition is a word-for-word repetition of [Mol02, Theorem 3.1] adapted to the setting of this paper.
Proposition A.4.
[Mol02, Theorem 3.1] We have
@@
Denote by $\alpha=N+1$. Recall the trick of Example˜3.6. Denote by $T_{1\ldots N}^{1\ldots j\mapsto\alpha\ldots N}(u)$ the quantum minor with values in $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N+1})$ corresponding to the indices
$(b_{1},\ldots,b_{N})=(1,\ldots,N),\qquad(a_{1},\ldots,a_{N})=(1,\ldots,\alpha,\ldots,N),$
in the notations of (1.9), where $\alpha$ is on the $j$-th place. By Proposition˜1.1 and (1.12), we have
$T_{1\ldots N}^{1\ldots j\mapsto\alpha\ldots N}(u)=\sum_{i=1}^{N}(-1)^{i+j}E_{\alpha i}T_{1\ldots\hat{i}\ldots N}^{1\ldots\hat{j}\ldots N}(u)=\sum_{i=1}^{N}\phi_{i}\hat{T}_{ij}(u).$
@@
$PT_{1\ldots N}^{1\ldots j\mapsto\alpha\ldots N}(-u+N\hbar-\hbar)=\sum_{I}(-1)^{s}\prod_{k\in\bar{I}}(-u+w_{k})P\phi_{i}E_{ii_{2}}\ldots E_{i_{s}j}\cdot\prod_{l>j}(-u+w_{l}),$
where $I=\{1\leq i=i_{1}<i_{2}\ldots<i_{s}<j\}$ and $\bar{I}$ is the complement of $I$ in the set $\{1,\ldots,j-1\}$. Using properties (4.6) again and pushing terms of $\mathfrak{n}_{+}$ to the right, we get
$(-1)^{N+1}\sum_{j=1}^{N}PT_{1\ldots N}^{1\ldots j\mapsto\alpha\ldots N}(-u+N\hbar-\hbar)v_{j}P=\sum_{i=1}^{N}\prod_{k<i}(u-w_{k})\cdot\prod_{k>i}(u-w_{k}+\hbar)P\phi_{i}v_{i}P,$

### v4

@@ 8.2 Monopole operators
$S^{k}\mathbf{C}^{N}\ni v_{1}^{\otimes k}\mapsto\Biggl(\sum_{i=1}^{N}\prod_{j\neq i}\frac{1}{w_{i}-w_{j}}\mathbf{u}_{i}^{-1}\Biggr)^{k}=:\sum_{J}c_{J}(w)\mathbf{u}_{j_{1}}^{-1}\cdots\mathbf{u}_{j_{k}}^{-1},$
where $J$ runs over all multisubsets of $\{1,\dots,N\}$ and $c_{J}(w)\in\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}$ are some functions. In what follows, we will use Proposition 5.12 and notations thereof. By Proposition 8.17, we need to compute
$\sum_{J}c_{J}(w)f\bigl(\vec{Y}\bigr)\mathbf{u}^{-1}_{j_{1}}\cdots\mathbf{u}^{-1}_{j_{k}},\qquad f\bigl(\vec{Y}\bigr)=\frac{1}{k!}\prod_{i\neq j}(Y_{i}-Y_{j}-\hbar).$
@@
Therefore, the only non-zero indices in the sum are permutations of $\{1,\dots,k\}$. Moreover, by $W$-invariance of Remark 8.11, it is enough to compute the coefficient $c_{\{1,\dots,k\}}(w)$. By an easy induction on $k$, we have
$c_{\{1,\dots,k\}}(w)=k!\prod_{1\leq i\neq j\leq k}\frac{1}{w_{i}-w_{j}-\hbar}\cdot\prod_{\begin{subarray}{c}i=1,\dots,k,\\ j=k+1,\dots,N\end{subarray}}\frac{1}{w_{i}-w_{j}}\mathbf{u}^{-1}_{1}\cdots\mathbf{u}^{-1}_{k}.$
By Proposition 5.12,
$c_{\{1,\dots,k\}}(w)f\bigl(\vec{Y}\bigr)\mathbf{u}_{1}^{-1}\cdots\mathbf{u}_{k}^{-1}=\prod_{\begin{subarray}{c}i=1,\dots,k,\\ j=k+1,\dots,N\end{subarray}}\frac{1}{w_{i}-w_{j}}\mathbf{u}_{1}^{-1}\cdots\mathbf{u}_{k}^{-1}.$
@@
$\bigl\langle g\cdot v_{\omega_{k}},v^{*}_{\omega_{k}}\bigr\rangle\mapsto\sum_{\begin{subarray}{c}I\subset\{1,\dots,N\},\\ |I|=k\}\end{subarray}}\prod_{\begin{subarray}{c}i\in I,\\ j\in\bar{I}\end{subarray}}\frac{1}{w_{i}-w_{j}}\prod_{i\in I}\mathbf{u}_{i}^{-1},$
@@
This is a rational analog of the simplest operator in the M-system in its shuffle realization, see [46, formula (104)] and [46, Proposition 6.13]. By [16, Remark 9.19], its operators correspond to the $K$-theory classes of line bundles $\mathcal{O}(-k)$ restricted to the orbits $\mathrm{Gr}^{\omega_{i}}$ as in Section 8.1, where $\mathcal{O}(1)$ is the theta bundle on the affine Grassmannian. Therefore, Proposition 8.17 and Theorem 8.18 are in accordance with the derived Satake isomorphism of Theorem 8.15, since by [14, Lemma 2.6], we have $\bigl\langle g\cdot v_{\omega_{i}},v_{\omega_{i}}^{*}\bigr\rangle\mapsto[\mathrm{Gr}^{\omega_{i}}]$, up to a scalar factor. It would be interesting to find a rational analog of the full $M$-system.
@@ Appendix A Technical lemmas
The following result is used in the proof of Proposition 3.3. Recall that
@@
$\prod_{i=1}^{l}G^{\theta}(u-z_{i},\kappa)=\sum_{i=1}^{l}G^{\theta}(u-z_{i},l\kappa)\cdot\prod_{\begin{subarray}{c}j=1,\dots,l,\\ j\neq i\end{subarray}}G^{\theta}(z_{i}-z_{j},\kappa).$
@@
$\frac{\theta(u-z_{i}+l\kappa)\theta(u-z_{l+1}+\hbar)}{\theta(u-z_{i})\theta(l\kappa)\theta(u-z_{l+1})\theta(\kappa)}$
$\qquad{}=\frac{\theta(u-z_{l+1}+l\kappa+\kappa)\theta(z_{l+1}-z_{i}+l\kappa)}{\theta(u-z_{l+1})\theta(l\kappa+\kappa)\theta(z_{l+1}-z_{i})\theta(l\kappa)}+\frac{\theta(z_{i}-z_{l+1}+\kappa)\theta(u-z_{i}+l\kappa+\kappa)}{\theta(z_{i}-z_{l+1})\theta(\kappa)\theta(u-z_{i})\theta(l\kappa+\kappa)}$
by multiplying by common denominator and taking
@@
The following result about symmetric polynomials is used in Proposition 6.7. Define symmetric polynomials $p^{\hbar}_{a}\bigl(\vec{\Omega}\bigr)\in\mathbf{C}[\hbar][\Omega_{1},\dots,\Omega_{k}]^{S_{k}}$ by
$\prod_{i=1}^{k}\frac{u-\Omega_{i}+\hbar}{u-\Omega_{i}}=1+\hbar\sum_{a=0}^{\infty}p_{a}^{\hbar}\bigl(\vec{\Omega}\bigr)u^{-a-1}\in\mathbf{C}[\hbar][\Omega_{1},\dots,\Omega_{k}]^{S_{k}}\bigl[\hskip-0.56905pt\bigl[u^{-1}\bigr]\hskip-0.56905pt\bigr],$
where we understand each term in the product as in Remark 5.8.
@@
    $p_{a}^{\hbar}\bigl(\vec{\Omega}\bigr)=\sum_{i=1}^{k}\Omega_{i}^{a}\prod_{j\neq i}\frac{\Omega_{i}-\Omega_{j}+\hbar}{\Omega_{i}-\Omega_{j}}.$ (A.1)
    In particular, $p^{0}_{a}\bigl(\vec{\Omega}\bigr)=\sum_{i=1}^{k}\Omega_{i}^{a}$ is the standard power sum symmetric function.
  2. $2.$
    The functions $p^{\hbar}_{a}\bigl(\vec{\Omega}\bigr)$ for $1\leq a\leq k$ freely generate the algebra of symmetric functions, in other words, there is an isomorphism
    $\mathbf{C}[\hbar][\Omega_{1},\dots,\Omega_{k}]^{S_{k}}\cong\mathbf{C}[\hbar]\bigl[p^{\hbar}_{1}\bigl(\vec{\Omega}\bigr),\dots,p^{\hbar}_{k}\bigl(\vec{\Omega}\bigr)\bigr].$
@@
$\mathbf{C}[\hbar][\Omega_{1},\dots,\Omega_{k}]^{S_{k}}\cong\mathbf{C}[\hbar]\bigl[p^{0}_{1}\bigl(\vec{\Omega}\bigr),\dots,p^{0}_{k}\bigl(\vec{\Omega}\bigr)\bigr],$
since $p^{0}_{a}\bigl(\vec{\Omega}\bigr)=\sum_{i=1}^{k}\Omega_{i}^{a}$. Let us consider the right-hand side as the coordinate ring of the affine space $\mathbb{A}^{k}\times\mathbb{A}^{1}$, where $\mathbb{A}^{1}$ corresponds to $\hbar$. There is a $\mathbf{C}^{*}$-action on $\mathbf{C}[\hbar][\Omega_{1},\dots,\Omega_{k}]$ such that all the variables $\Omega_{i}$ and $\hbar$ have weight 1. It induces a $\mathbf{C}^{*}$-action on $\mathbf{C}[\hbar][\Omega_{1},\dots,\Omega_{k}]^{S_{k}}$ with positive weights; in particular, it contracts $\mathbb{A}^{k}\times\mathbb{A}^{1}$ to the origin. In fact, under the isomorphism above, this action is linear and diagonal with weights $(1,2,\dots,k,1)$.
Consider the map $\varphi_{\hbar}\colon\mathbb{A}^{k}\times\mathbb{A}^{1}\rightarrow\mathbb{A}^{k}\times\mathbb{A}^{1}$ defined on the ring of functions by $p^{0}_{a}\bigl(\vec{\Omega}\bigr)\mapsto p^{\hbar}_{a}\bigl(\vec{\Omega}\bigr)$ for every $a$ and $\hbar\mapsto\hbar$. It follows from (A.1) that this map is $\mathbf{C}^{*}$-equivariant. Moreover, it is clear from weight considerations and explicit calculations for $a=1$ that the linear part of $\varphi_{\hbar}$ is
$\mathrm{d}\varphi_{\hbar}\bigl(p^{0}_{a}\bigl(\vec{\Omega}\bigr)\bigr)=p_{a}^{0}\bigl(\vec{\Omega}\bigr)+\delta_{a1}\frac{k(k-1)}{2}\hbar,\qquad\mathrm{d}\varphi_{\hbar}(\hbar)=\hbar.$
Therefore, the differential of $\varphi_{\hbar}$ at zero is invertible. By the proof of [19, Lemma 2.1], it is an isomorphism. In particular, the polynomials $\bigl\{p^{\hbar}_{a}\bigl(\vec{\Omega}\bigr)\bigr\}$ generate the ring of symmetric functions. ∎
In what follows, we use the notations of Section 5 for $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes\bigl(S^{\bullet}\mathbf{C}^{N}\otimes S^{\bullet}\bigl(\mathbf{C}^{N}\bigr)^{*}\bigr)$ (more precisely, for its Mickelsson algebra).
@@
$P\sum_{1\leq i_{1}<\dots<i_{s}<i}(w_{i}-w_{i_{1}})^{-1}\cdots(w_{i}-w_{i_{s}})^{-1}\phi_{i_{1}}E_{i_{1}i_{2}}\cdots E_{i_{s}i}\cdot v_{j}P$
$\qquad{}=Pv_{j}\sum_{1\leq i_{1}<\dots<i_{s}<i}(w_{i}-w_{i_{1}})^{-1}\cdots(w_{i}-w_{i_{s}})^{-1}\phi_{i_{1}}E_{i_{1}i_{2}}\cdots E_{i_{s}i}P=Pv_{j}\phi_{i}P,$
@@
$P\sum_{N\geq j_{1}>\dots>j_{s}>j}(w_{j}-w_{j_{1}})^{-1}\cdots(w_{i}-w_{j_{s}})^{-1}v_{j_{1}}E_{j_{2}j_{1}}\cdots E_{jj_{s}}E_{i_{1}i_{2}}\cdot\phi_{i}P=P\phi_{i}v_{j}P.$
@@
$\varpi_{N}\colon\ \mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes\bigl(S^{\bullet}\mathbf{C}^{N}\otimes S^{\bullet}\bigl(\mathbf{C}^{N}\bigr)^{*}\bigr)\rightarrow\bigl[\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes\bigl(S^{\bullet}\mathbf{C}^{N}\otimes S^{\bullet}\bigl(\mathbf{C}^{N}\bigr)^{*}\bigr)\bigr]^{\mathrm{opp}},$
$\varpi_{N}(v_{i})=\phi_{i},\qquad\varpi_{N}(\phi_{i})=v_{i},$
extending the one (2.5) of $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})$. By Remark 5.3, we have $\varpi_{N}(P)=P$, therefore,
@@
The proof of the following proposition is a word-for-word repetition of [37, Theorem 3.1] adapted to the setting of this paper.
Proposition A.4 ([37, Theorem 3.1]).
We have
@@
Denote by $\alpha=N+1$. Recall the trick of Example 4.6. Denote by $T_{1\dots N}^{1\dots j\mapsto\alpha\dots N}(u)$ the quantum minor with values in $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N+1})$ corresponding to the indices
$(b_{1},\dots,b_{N})=(1,\dots,N),\qquad(a_{1},\dots,a_{N})=(1,\dots,\alpha,\dots,N),$
in the notations of (2.7), where $\alpha$ is on the $j$-th place. By Proposition 2.1 and (2.9), we have
$T_{1\dots N}^{1\dots j\mapsto\alpha\dots N}(u)=\sum_{i=1}^{N}(-1)^{i+j}E_{\alpha i}T_{1\dots\hat{i}\dots N}^{1\dots\hat{j}\dots N}(u)=\sum_{i=1}^{N}\phi_{i}\hat{T}_{ij}(u).$
@@
$PT_{1\dots N}^{1\dots j\mapsto\alpha\dots N}(-u+N\hbar-\hbar)=\sum_{I}(-1)^{s}\prod_{k\in\bar{I}}(-u+w_{k})P\phi_{i}E_{ii_{2}}\cdots E_{i_{s}j}\cdot\prod_{l>j}(-u+w_{l}),$
where $I=\{1\leq i=i_{1}<i_{2}<\dots<i_{s}<j\}$ and $\bar{I}$ is the complement of $I$ in the set $\{1,\dots,j-1\}$. Using properties (5.5) again and pushing terms of $\mathfrak{n}_{+}$ to the right, we get
$(-1)^{N+1}\sum_{j=1}^{N}PT_{1\dots N}^{1\dots j\mapsto\alpha\dots N}(-u+N\hbar-\hbar)v_{j}P=\sum_{i=1}^{N}\prod_{k<i}(u-w_{k})\cdot\prod_{k>i}(u-w_{k}+\hbar)P\phi_{i}v_{i}P,$